the perimeter of a rectangle is 30 inches, and its area is 56 square inches .find the length and width?
Answer by hheuler(6)

(Eq 1) P = 2L + 2W = 30
(Eq 2) A = LW = 56
From (Eq 1): L + W = 15 or L = 15 - W (Eq 3)
Substitute (Eq 3) into (Eq 2): (15 - W)W = 56 = 8(7)= (15 - 7)7 (Eq 4)
So the width is 7 by inspection of (Eq 4)and from (Eq 1), the length is 8.
QED